The number of people arriving at a post office in a small town follows a Poisson distribution. On average, five people arrive at the post office in one minute. Find the probability that in a given minute, 

A) more than two people arrive at the post office

B) less than eleven people arrive at the post office

c) between six and ten people arrive at the post office


1
Expert's answer
2021-06-09T07:33:10-0400

Let "X=" the number of people arriving at a post office: "X\sim Po(\lambda)."


"P(X=x)=\dfrac{e^{-\lambda}\cdot\lambda^x}{x!}"

Given "\lambda=5"

A)


"P(X>2)=1-P(X=0)-P(X=1)-P(X=2)"

"=1-\dfrac{e^{-5}\cdot5^0}{0!}-\dfrac{e^{-5}\cdot5^1}{1!}-\dfrac{e^{-5}\cdot5^2}{2!}"

"\approx0.875348"

B)


"P(X<11)=P(X=0)-P(X=1)-P(X=2)"

"+P(X=3)+P(X=4)+P(X=5)"

"+P(X=6)+P(X=7)+P(X=8)"

"+P(X=9)+P(X=10)"

"\approx0.986305"

C)


"P(6\leq X\leq10)=P(X=6)+P(X=7)"

"+P(X=8)+P(X=9)+P(X=10)"

"\approx0.370344"